OpenStudy (anonymous):

Calculate the molality of 2.0 mol HCl in 175 g water

OpenStudy (anonymous):

mole = mass / m.mass mass= mole X m.mass = 2 X 36.5 = 73 g of HCl its mean 73 g of HCl are dissolved in 175 g. 175 g of water contain = 73 g 1000 g of water contain = 73/175 X 1000 = 417 g/kg